# A Tri-Modal Dataset and a Baseline System for Tracking Unmanned Aerial Vehicles

### v1

@@ I INTRODUCTION
The rapid proliferation of unmanned aerial vehicles (UAVs) has significantly impacted various sectors, including civilian logistics, aerial cinematography, infrastructure inspection, emergency response, and environmental monitoring [application1, application2, application3]. Their adaptability and operational efficiency have rendered them indispensable across both commercial and industrial domains. However, this expansive use introduces critical security vulnerabilities, raising substantial concerns regarding personal privacy, public safety, and national security. Consequently, the development of robust UAV detection and tracking systems has emerged as an urgent technological imperative [DUT-Anti-UAV].
@@
As a fundamental component of low-altitude security systems, UAV tracking demands increasingly robust algorithms to counter evolving and sophisticated UAV threats. The core objective is to achieve continuous, real-time, and stable trajectory estimation of UAVs across diverse and challenging environments. In contrast to traditional radar [radar1] or radio frequency [RF] based countermeasures, vision-based UAV tracking has gained prominence, largely driven by significant advances in computer vision [YOLOX, detr] and visual tracking algorithms [DeepSORT, bytetrack, botsort, trackformer].
@@
To address these challenges, a growing body of research [Anti-UAV, Anti-UAV410, Anti-UAV600, 2th-antiuav, 3th-antiuav, SiamSTA, GLTF-MA, UTTracker] has been dedicated to developing visual object tracking methods tailored for UAV-targeted applications. Nevertheless, the majority of these efforts remain confined to single visual modalities and single-object tracking (SOT) frameworks. Although multi-UAV tracking has recently attracted attention [MOT-FLY, Dist-Tracker, strong-baseline], existing approaches still rely predominantly on a single type of sensor data. Single-modality tracking is inherently limited by sensor-specific constraints, leading to notable performance bottlenecks in practical scenarios. For instance, RGB cameras, which operate in the visible spectrum, are highly susceptible to performance degradation under poor illumination conditions, such as at night, in heavy fog, or during rain, leading to significant noise and loss of detail. In contrast, infrared (IR) sensors detect thermal radiation and can operate in darkness, but are often susceptible to false positives caused by birds or other heat-emitting objects, while lacking the ability to discern textual or fine-grained visual features of UAVs. These limitations substantially reduce the operational effectiveness of single-modality trackers in real-world anti-UAV operations.
@@
Similarly, SOT paradigms, which typically require manual initialisation of a single target in the first frame, preserving tracking of individual objects but are ill-suited for scenarios involving variable numbers of UAVs. In realistic UAV missions, threats may emerge, disappear, or re-enter the field of view abruptly and without prior knowledge, making manual initialisation impractical. Although methods such as [Anti-UAV600] attempt to reduce reliance on manual prior information, they remain limited to single-modality SOT tracking, thus cannot handle multi-object scenarios. Therefore, multi-modal multi-object tracking (MM-MOT) presents a promising direction by leveraging complementary sensor information to enhance robustness across diverse conditions, while also supporting autonomous detection and tracking of multiple uncertain UAV targets. This capability is critical for responsive and scalable anti-UAV systems in dynamic and complex environments.
@@
FL-Drones [FL-Drones] Det RGB 14 39K 640×480-752×480 2016 TPAMI
NPS-Drones [NPS-Drones] Det RGB 50 70K 1280×760-1920×1280 2016 IROS
Det-Fly [Det-Fly] Det RGB 13K 3840×2160 2021 RA-L
Drone-vs-Bird [Drone-vs-Bird] Det RGB 77 105K 720×576-3840×2160 34×23 2021 AVSS
UETT4K Anti-UAV [UETT4K-Anti-UAV] Det RGB 18 33K 3840×2160 2025 IEEE Access
USC Drone [USC-Drone] SOT RGB 60 20K 1920×1080 114×83 2017 APSIPA ASC
Anti-UAV [Anti-UAV] SOT RGB, IR 348 297K 1920×1080 (RGB) 640×512 (IR) 125×63 (RGB) 52×30 (IR) 2021 TMM
Anti-UAV600 [Anti-UAV600] SOT IR 600 723K 640×512 30×19 2023 CoRR
Anti-UAV410 [Anti-UAV410] SOT IR 410 438K 640×512 2024 TPAMI
Halmstad Drone [Halmstad-Drone] Det/SOT RGB, IR 650 203K 640×512 (RGB) 320×256 (IR) 2021 ICPR
DUT Anti-UAV [DUT-Anti-UAV] Det/SOT RGB 20 35K 160×240-3744×5616 70×31 2022 T-ITS
UAVSwarm [UAVSwarm] MOT RGB 72 12K 446×276-1919×1079 2022 Remote Sensing
MOT-Fly [MOT-FLY] MOT RGB 16 11K 1920×1080 2023 ICUS
4th Anti-UAV challenge [4th-Anti-UAV-workshop] MOT IR 300 225K 640×512 2025 CVPR workshop
@@
The name of anti-UAV qualifying these datasets has understandably been coined in [DUT-Anti-UAV, Anti-UAV, Anti-UAV410, Anti-UAV600, 2th-antiuav, 3th-antiuav, SiamSTA, GLTF-MA] to emphasise the defensive purpose of UAV tracking. However, we shall not perpetuate the use of this term as it is ambiguous and counterintuitive. The task is to track simultaneously several very challenging objects, and fundamentally, it is a multi-object tracking problem. If tracked successfully, an appropriate action can be taken to neutralise any threat posed by the objects. However, the discussion of any such action is beyond the scope of this paper. We focus simply on the topic of multi-UAV tracking, and its peculiarities, to advance multi-object tracking in general, and multi-UAV tracking in particular. In this context, we have collected the first large-scale tri-modal UAV tracking dataset. It can be used for multimodal tracking per se, but it can as well be used for related research topics, such as multimodal UAV detection, for studies of adversarial attacks on tracking systems, and for adversarial attack detection.
@@
In general, multi-object tracking (MOT), also referred to as multi-target tracking, aims to detect and localise multiple targets across video sequences while maintaining consistent identities over time [DeepSORT]. Compared to conventional MOT tasks such as pedestrian tracking [mot15, mot20], multi-UAV tracking presents significantly greater challenges in identity preservation [Dist-Tracker, strong-baseline]. These difficulties stem from three primary factors. First, UAVs often exhibit fast and irregular motion patterns, causing conventional Kalman filters [KF] to accumulate large prediction errors, which leads to identity mismatches in IoU-based association modules. Second, unlike pedestrians that often possess distinctive appearance features, UAVs typically share highly similar or identical visual characteristics, which greatly limits the discriminative power of appearance-based Re-ID models. Third, UAVs typically appear as small objects with limited texture details, making them prone to being confused with background clutter, which results in elevated false negatives and false positives.
@@
Due to these challenges, conventional MOT frameworks [DeepSORT, bytetrack, botsort, boostTrack, boostTrack++] that rely primarily on Kalman filtering and appearance embeddings exhibit considerable limitations in multi-UAV scenarios. In response, the emerging event modality visual sensing technology, capturing asynchronous, sparse motion streams with high temporal resolution [ceutrack, fe108, eventvot, TENet], has recently gained increasing attention in visual tracking. By directly encoding intensity changes over time, event cameras enable robust motion-based reasoning that helps disambiguate small, fast-moving objects such as UAVs. This modality provides strong motion cues that are complementary to conventional RGB or IR sensors, thereby reducing identity switches caused by appearance ambiguity or erratic motion patterns.
@@
Given the current absence of multi-modal frameworks in existing multi-UAV tracking studies [MOT-FLY, strong-baseline, Dist-Tracker], we propose a novel multi-modal multi-object tracking approach designed specifically to address the identity discrimination challenges in UAV scenarios. Our method, termed Multi-Modal-Aware Simple Online and Realtime Tracking (MMA-SORT), effectively integrates RGB, IR, and event modalities within a unified tracking framework. Beyond conventional motion prediction using Kalman filters and appearance modelling via Re-ID features, we introduce an additional motion embedding derived directly from event data. This embedding captures fine-grained temporal dynamics that are complementary to appearance cues, significantly enhancing the discrimination of UAV targets with similar appearance and reducing identity switches caused by background clutter or erratic motion. Notably, the proposed motion embedding is extracted in a training-free manner directly from raw event streams, avoiding the computational overhead associated with deep feature extraction, thereby guaranteeing high inference speed. The MMA-SORT framework effectively combines the strengths of three complementary sensing modalities, offering robust performance in challenging conditions where single-modality or conventional dual-modality trackers tend to fail.
@@
Furthermore, in multi-modal visual tasks that demand high localisation accuracy, spatio-temporal misalignment across modalities presents a critical challenge. Such misalignment arises from differences in imaging mechanisms, field of view, data transmission latency, and resolution among heterogeneous sensors [CSOM, align2]. While often tolerable for larger objects, even slight misalignments cause severe localisation errors for small UAV targets, which may span only a few pixels, significantly undermining the effectiveness of multi-modal fusion. To tackle this issue, we propose an Offset-Guided Adaptive Alignment Module (OGAA), which performs feature-level alignment across modalities before fusion. OGAA employs two complementary strategies, i.e., deformable convolution [def-conv] and spatial transformer networks [STN], to dynamically adjust feature maps and reduce spatial discrepancies. Following alignment, a lightweight adaptive dynamic fusion module is introduced to balance the contributions of different modalities, further enhancing fusion quality and robustness. Together, these components form a cohesive fusion pipeline that significantly improves tracking accuracy for small and fast-moving UAV targets.
@@
Existing benchmark datasets for visual UAV tracking and detection tasks, such as those summarised in Table I, focus on single-modal detection or single-object tracking [Anti-UAV, Anti-UAV410, Anti-UAV600]. Although a few datasets, including AntiUAV and Halmstad Drone, incorporate multiple modalities (e.g., RGB and infrared), they remain confined to single-object tracking scenarios. As a result, there is a conspicuous absence of large-scale, publicly available datasets supporting multi-modal multi-object tracking for UAVs, hindering the development of robust solutions. To bridge this gap, we introduce MM-UAV, the first large-scale Multi-Modal UAV Tracking dataset, designed specifically for multi-object tracking across three sensing modalities, i.e., RGB, infrared, and event data. Comprising 1,321 sequences (1,200 for training and 121 for testing), with approximately 2.8 million frames per modality, MM-UAV offers unprecedented scale and diversity. Notably, we adopt a stricter trajectory annotation protocol, i.e., UAVs that leave and re-enter the scene retain their original identity, avoiding identity fragmentation common in other datasets. The key characteristics of MM-UAV can be summarised as: (1) smaller target sizes, (2) larger data scale, (3) richer modality variety, and (4) more consistent identity annotations. To the best of our knowledge, it is the largest and most comprehensive dataset of its kind to date. A detailed description and statistical analysis of the dataset are provided in Sec. III.
In summary, this work presents the following key contributions:
@@ II-A UAV-Targeted Tracking Solutions
Unlike UAV-view tracking [fan2020visdrone], where the UAV serves as an aerial observer to perform air-to-ground tracking of objects such as pedestrians and vehicles within ground scenes, UAV-targeted tracking considers UAVs as objects of interest, aiming to achieve continuous and stable trajectory tracking of UAVs in complex low-altitude scenarios. Compared to general single-object tracking tasks, which rely on distinctive appearance cues and moderate scale variations, UAV targets are typically small, visually indistinct, fast-moving, and often concealed. As a result, in UAV-targeted scenarios, single-object tracking is required to maintain the trajectory of an initialised target across video frames as well as to overcome additional challenges such as weak appearance discriminability, frequent disappearance, and intermittent reappearance. Most existing solutions are based on the Siamese network architecture, achieving target localisation through cross-frame template feature matching. Representative examples include early CNN-based solutions like SiamFC [siamfc], SiamRPN [siamRPN], and SiamRPN++ [siamRPN++], as well as Transformer-based trackers such as Stark [stark], OSTrack [ostrack], and MixFormer [mixformer]. However, these trackers are optimised primarily on large-scale general tracking datasets, such as GOT-10k [got10k], TrackingNet [trackingnet], and LaSOT [lasot], lacking specific descriptions for UAVs.
@@
Given the specific appearance and the long-term nature of the UAV tracking tasks, many anti-UAV SOT solutions employ a “local tracking + global re-detection” framework. For instance, SiamSTA [SiamSTA] leverages a spatio-temporal attention mechanism and a three-stage global re-detection strategy to track fast-moving UAVs. UTTracker [UTTracker] performs multi-region local tracking with temporal cues, combined with global detection and background correction modules, to handle drastic scale variations, frequent disappearance, and camera motion. Despite their achievements, these solutions are limited to single-object tracking and cannot handle complex scenarios involving collaborative intrusion by unknown numbers of UAVs. Thus, multi-UAV tracking has become a focus to meet practical demands.
@@
In general, Multi-Object Tracking (MOT) focuses on detecting and tracking multiple targets across frames. Building on this paradigm, multi-UAV tracking aims to associate trajectories of multiple UAVs without prior target initialisation, representing a crucial step toward practical anti-UAV technologies. Current mainstream MOT frameworks fall into two paradigms: tracking-by-detection (e.g., DeepSORT [DeepSORT], ByteTrack [bytetrack], BoTSORT [botsort], BoostTrack [boostTrack]) and tracking-by-query (e.g., TransTrack [transtrack], Trackformer [trackformer], MOTR [motr], MeMOTR [memotr]). However, UAV-specific challenges significantly degrade their performance. First, high appearance similarity reduces the effectiveness of traditional ReID-based embeddings. Second, fast and irregular motion violates linear motion assumptions in Kalman filtering, leading to prediction errors and ID switches. Third, the small UAV target size increases detection failures, further complicating data association.
@@
These limitations highlight the need for UAV-oriented solutions in multi-object tracking paradigms. Drawing on this, recent studies have explored adapting generic MOT pipelines to small targets, particularly in thermal infrared videos. Dist-Tracker [Dist-Tracker] combines a scale-shape-quality detector based on YOLOv12 [yolov12] with a fusion of L2-IoU tracker to handle low contrast, scale variations, and motion instability. Strong Baseline [strong-baseline] achieves competitive results using the YOLOv12 and BoT-SORT framework with customised training/inference strategies. Nevertheless, the progress in multi-UAV tracking has been limited, with most approaches confined to single-modality inputs, resulting in insufficient robustness under challenging conditions.

### v2

@@ I INTRODUCTION
The rapid proliferation of unmanned aerial vehicles (UAVs) has significantly impacted various sectors, including civilian logistics, aerial cinematography, infrastructure inspection, emergency response, and environmental monitoring. Their adaptability and operational efficiency have rendered them indispensable across both commercial and industrial domains. However, this expansive use introduces critical security vulnerabilities, raising substantial concerns regarding personal privacy, public safety, and national security. Consequently, the development of robust UAV detection and tracking systems has emerged as an urgent technological imperative [66].
@@
As a fundamental component of low-altitude security systems, UAV tracking demands increasingly robust algorithms to counter evolving and sophisticated UAV threats. The core objective is to achieve continuous, real-time, and stable trajectory estimation of UAVs across diverse and challenging environments. In contrast to traditional radar [67] or radio frequency [3] based countermeasures, vision-based UAV tracking has gained prominence, largely driven by significant advances in computer vision [23, 7, 13] and visual tracking algorithms [40, 62, 2, 38, 54, 30].
@@
To address these challenges, a growing body of research [28, 25, 70, 65, 64, 24, 35, 57] has been dedicated to developing visual object tracking methods tailored for UAV-targeted applications. Nevertheless, the majority of these efforts remain confined to single visual modalities and single-object tracking (SOT) frameworks. Although multi-UAV tracking has recently attracted attention [14, 50, 11], existing approaches still rely predominantly on a single type of sensor data. Single-modality tracking is inherently limited by sensor-specific constraints, leading to notable performance bottlenecks in practical scenarios. For instance, RGB cameras, which operate in the visible spectrum, are highly susceptible to performance degradation under poor illumination conditions, such as at night, in heavy fog, or during rain, leading to significant noise and loss of detail. In contrast, infrared (IR) sensors detect thermal radiation and can operate in darkness, but are often susceptible to false positives caused by birds or other heat-emitting objects, while lacking the ability to discern textual or fine-grained visual features of UAVs. These limitations substantially reduce the operational effectiveness of single-modality trackers in real-world anti-UAV operations.
@@
Similarly, SOT paradigms, which typically require manual initialisation of a single target in the first frame, preserving tracking of individual objects but are ill-suited for scenarios involving variable numbers of UAVs. In realistic UAV missions, threats may emerge, disappear, or re-enter the field of view abruptly and without prior knowledge, making manual initialisation impractical. Although methods such as [70] attempt to reduce reliance on manual prior information, they remain limited to single-modality SOT tracking, thus cannot handle multi-object scenarios. Therefore, multi-modal multi-object tracking (MM-MOT) presents a promising direction by leveraging complementary sensor information to enhance robustness across diverse conditions, while also supporting autonomous detection and tracking of multiple uncertain UAV targets. This capability is critical for responsive and scalable anti-UAV systems in dynamic and complex environments.
@@
FL-Drones [41] Det RGB 14 39K 640×480-752×480 2016 TPAMI
NPS-Drones [34] Det RGB 50 70K 1280×760-1920×1280 2016 IROS
Det-Fly [68] Det RGB 13K 3840×2160 2021 RA-L
Drone-vs-Bird [15] Det RGB 77 105K 720×576-3840×2160 34×23 2021 AVSS
UETT4K Anti-UAV [4] Det RGB 18 33K 3840×2160 2025 IEEE Access
USC Drone [12] SOT RGB 60 20K 1920×1080 114×83 2017 APSIPA ASC
Anti-UAV [28] SOT RGB, IR 348 297K 1920×1080 (RGB) 640×512 (IR) 125×63 (RGB) 52×30 (IR) 2021 TMM
Anti-UAV600 [70] SOT IR 600 723K 640×512 30×19 2023 CoRR
Anti-UAV410 [25] SOT IR 410 438K 640×512 2024 TPAMI
Halmstad Drone [46] Det/SOT RGB, IR 650 203K 640×512 (RGB) 320×256 (IR) 2021 ICPR
DUT Anti-UAV [66] Det/SOT RGB 20 35K 160×240-3744×5616 70×31 2022 T-ITS
UAVSwarm [49] MOT RGB 72 12K 446×276-1919×1079 2022 Remote Sensing
MOT-Fly [14] MOT RGB 16 11K 1920×1080 2023 ICUS
4th Anti-UAV challenge [1] MOT IR 300 225K 640×512 2025 CVPR workshop
@@
The name of anti-UAV qualifying these datasets has understandably been coined in [66, 28, 25, 70, 65, 64, 24, 35] to emphasise the defensive purpose of UAV tracking. However, we shall not perpetuate the use of this term as it is ambiguous and counterintuitive. The task is to track simultaneously several very challenging objects, and fundamentally, it is a multi-object tracking problem. If tracked successfully, an appropriate action can be taken to neutralise any threat posed by the objects. However, the discussion of any such action is beyond the scope of this paper. We focus simply on the topic of multi-UAV tracking, and its peculiarities, to advance multi-object tracking in general, and multi-UAV tracking in particular. In this context, we have collected the first large-scale tri-modal UAV tracking dataset. It can be used for multimodal tracking per se, but it can as well be used for related research topics, such as multimodal UAV detection, for studies of adversarial attacks on tracking systems, and for adversarial attack detection.
@@
In general, multi-object tracking (MOT), aims to detect and localise multiple targets across video sequences while maintaining consistent identities over time [40]. Compared to conventional MOT tasks such as pedestrian tracking [31, 18], multi-UAV tracking presents significantly greater challenges in identity preservation [50, 11]. These difficulties stem from three primary factors. First, UAVs often exhibit fast and irregular motion patterns, causing conventional Kalman filters [29] to accumulate large prediction errors, which leads to identity mismatches in IoU-based association modules. Second, unlike pedestrians that often possess distinctive appearance features, UAVs typically share highly similar or identical visual characteristics, which greatly limits the discriminative power of appearance-based Re-ID models. Third, UAVs typically appear as small objects with limited texture details, making them prone to being confused with background clutter, which results in elevated false negatives and false positives.
@@
Due to these challenges, conventional MOT frameworks [40, 62, 2, 43, 44] that rely primarily on Kalman filtering and appearance embeddings exhibit considerable limitations in multi-UAV scenarios. In response, the emerging event modality visual sensing technology, capturing asynchronous, sparse motion streams with high temporal resolution [47, 60, 52, 42], has recently gained increasing attention in visual tracking. By directly encoding intensity changes over time, event cameras enable robust motion-based reasoning that helps disambiguate small, fast-moving objects such as UAVs. This modality provides strong motion cues that complement conventional RGB or IR sensors, thereby reducing identity switches caused by appearance ambiguity or erratic motion patterns.
@@
Given the current absence of multi-modal frameworks in existing multi-UAV tracking studies [14, 11, 50], we propose a novel multi-modal multi-object tracking approach designed specifically to address the identity discrimination challenges in UAV scenarios. Our method, termed Multi-Modal-Aware Simple Online and Realtime Tracking (MMA-SORT), effectively integrates RGB, IR, and event modalities within a unified tracking framework. Beyond conventional motion prediction using Kalman filters and appearance modelling via Re-ID features, we introduce an additional motion embedding derived directly from event data. This embedding captures fine-grained temporal dynamics that are complementary to appearance cues, significantly enhancing the discrimination of UAV targets with similar appearance and reducing identity switches caused by background clutter or erratic motion. Notably, the proposed motion embedding is extracted in a training-free manner directly from raw event streams, avoiding the computational overhead associated with deep feature extraction, thereby guaranteeing high inference speed. The MMA-SORT framework effectively combines the strengths of three complementary sensing modalities, offering robust performance in challenging conditions where single-modality or conventional dual-modality trackers tend to fail.
@@
Furthermore, in multi-modal visual tasks that demand high localisation accuracy, spatio-temporal misalignment across modalities presents a critical challenge. Such misalignment arises from differences in imaging mechanisms, field of view, data transmission latency, and resolution among heterogeneous sensors [9, 10]. While often tolerable for larger objects, even slight misalignments cause severe localisation errors for small UAV targets, which may span only a few pixels, significantly undermining the effectiveness of multi-modal fusion. To tackle this issue, we propose an Offset-Guided Adaptive Alignment Module (OGAA), which performs feature-level alignment across modalities before fusion. OGAA employs two complementary strategies, i.e., deformable convolution [17] and spatial transformer networks [27], to dynamically adjust feature maps and reduce spatial discrepancies. Following alignment, a lightweight adaptive dynamic fusion module is introduced to balance the contributions of different modalities, further enhancing fusion quality and robustness. Together, these components form a cohesive fusion pipeline that significantly improves tracking accuracy for small and fast-moving UAV targets.
@@
Existing benchmark datasets for visual UAV tracking and detection tasks, such as those summarised in Table I, focus on single-modal detection or single-object tracking [28, 25, 70]. Although a few datasets, including AntiUAV and Halmstad Drone, incorporate multiple modalities (e.g., RGB and infrared), they remain confined to single-object tracking scenarios. As a result, there is a conspicuous absence of large-scale, publicly available datasets supporting multi-modal multi-object tracking for UAVs, hindering the development of robust solutions. To bridge this gap, we introduce MM-UAV, the first large-scale Multi-Modal UAV Tracking dataset, designed specifically for multi-object tracking across three sensing modalities, i.e., RGB, infrared, and event data. Comprising 1,321 sequences (1,200 for training and 121 for testing), with approximately 2.8 million frames per modality, MM-UAV offers unprecedented scale and diversity. Notably, we adopt a stricter trajectory annotation protocol, i.e., UAVs that leave and re-enter the scene retain their original identity, avoiding identity fragmentation common in other datasets. The key characteristics of MM-UAV can be summarised as: (1) smaller target sizes, (2) larger data scale, (3) richer modality variety, and (4) more consistent identity annotations. To the best of our knowledge, it is the largest and most comprehensive dataset of its kind to date. A detailed description and statistical analysis of the dataset are provided in Sec. III. In summary, this work presents the following key contributions:
@@ II-A UAV-Targeted Tracking Solutions
Unlike UAV-view tracking [20], where the UAV serves as an aerial observer to perform air-to-ground tracking of objects such as pedestrians and vehicles within ground scenes, UAV-targeted tracking considers UAVs as objects of interest, aiming to achieve continuous and stable trajectory tracking of UAVs in complex low-altitude scenarios. Compared to general single-object tracking tasks, which rely on distinctive appearance cues and moderate scale variations, UAV targets are typically small, visually indistinct, fast-moving, and often concealed. As a result, in UAV-targeted scenarios, single-object tracking is required to maintain the trajectory of an initialised target across video frames as well as to overcome additional challenges such as weak appearance discriminability, frequent disappearance, and intermittent reappearance. Most existing solutions are based on the Siamese network architecture, achieving target localisation through cross-frame template feature matching [54]. Representative examples include early CNN-based solutions like SiamFC [8], SiamRPN [33], and SiamRPN++ [32], as well as Transformer-based trackers such as Stark [55], OSTrack [56], and MixFormer [16]. However, these trackers are optimised primarily on large-scale general tracking datasets, such as GOT-10k [26], TrackingNet [39], and LaSOT [19], lacking specific descriptions for UAVs.
@@
Given the specific appearance and the long-term nature of the UAV tracking tasks, many anti-UAV SOT solutions employ a “local tracking + global re-detection” framework. For instance, SiamSTA [24] leverages a spatio-temporal attention mechanism and a three-stage global re-detection strategy to track fast-moving UAVs. UTTracker [57] performs multi-region local tracking with temporal cues, combined with global detection and background correction modules, to handle drastic scale variations, frequent disappearance, and camera motion. Despite their achievements, these solutions are limited to single-object tracking and cannot handle complex scenarios involving collaborative intrusion by unknown numbers of UAVs. Thus, multi-UAV tracking has become a focus to meet practical demands.
@@
In general, Multi-Object Tracking focuses on detecting and tracking multiple targets across frames. Building on this paradigm, multi-UAV tracking aims to associate trajectories of multiple UAVs without prior target initialisation, representing a crucial step toward practical anti-UAV technologies. Current mainstream MOT frameworks fall into two paradigms: tracking-by-detection (e.g., DeepSORT [40], ByteTrack [62], BoTSORT [2], BoostTrack [43]) and tracking-by-query (e.g., TransTrack [45], Trackformer [38], MOTR [58], MeMOTR [22]). However, UAV-specific challenges significantly degrade their performance. First, high appearance similarity reduces the effectiveness of traditional ReID-based embeddings. Second, fast and irregular motion violates linear motion assumptions in Kalman filtering, leading to prediction errors and ID switches. Third, the small UAV target size increases detection failures, further complicating data association.
@@
These limitations highlight the need for UAV-oriented solutions in multi-object tracking paradigms. Drawing on this, recent studies have explored adapting generic MOT pipelines to small targets, particularly in thermal infrared videos. Dist-Tracker [50] combines a scale-shape-quality detector based on YOLOv12 [48] with a fusion of L2-IoU tracker to handle low contrast, scale variations, and motion instability. Strong Baseline [11] achieves competitive results using the YOLOv12 and BoT-SORT framework with customised training/inference strategies. Nevertheless, the progress in multi-UAV tracking has been limited, with most approaches confined to single-modality inputs, resulting in insufficient robustness under challenging conditions.